# Quantifying and Mitigating Unimodal Biases in Multimodal Large Language Models: A Causal Perspective

### v2

@@ 2 A Causal Framework
Figure 2: Causal graph of MLLM’s Prediction on VQA problems. We use the green subgraph $G_{h}$ to represent the desired causal mechanisms and compare it with the undesired effects of vision and language biases. We quantify the causal effects of each factor by performing controlled interventions of the images ($I$) and of the question ($Q$, $T$).
@@ 2.1 Problem Setup
We consider an entity-centric VQA problem, denoted by $M$, consisting of a question $Q$ and an image $I$. The image depicts a specific entity and the question is related to the entity. The question $Q:=(S,T)$ is composed of two distinct elements: the core semantic content ${S}$, which conveys the genuine intent of the question, and the textual surface form ${T}$, which is irrelevant to the question’s core meaning. The model’s final answer/prediction is represented by $A$. In this paper, we use lowercase letters to represent an instance of its corresponding capital variable.
@@ 2.2 Causal Graph of MLLMs’ Prediction
${s}=f_{\text{c}}({q}),\ \ g=f_{{s}}({i}),$ (1)
where the cognitive process $f_{\text{c}}$ is employed to decipher the core semantic meaning $s$ of the question $q$. Then, the function $f_{s}$ is utilized to relate $s$ to the image $i$, yielding the final result $g$. These mechanisms are depicted within the green subgraph $G_{h}$ of Figure 2.
Contrarily, the process of a model to tackle the same VQA problem $m$ involves the following causal mechanisms:
$a=f_{\text{b}}({i},{q}),\vspace{-5pt}$ (2)
where $f_{\text{b}}$ functions as a black box, leaving uncertainty regarding which aspects of $q$ are considered by the model and its interaction with the image $i$.
To conduct this analysis, we draw all possible causal mechanisms that might take place in the complete causal graph in Figure 2. Notable causal mechanisms include:
@@
The model may attend directly to the image $I$ via a causal path ${I}\rightarrow A$, leading to the emergence of vision bias, e.g., the focus on the image in Figure 1 (b).
Language Bias
The model may directly process the question $Q$ in two ways: by focusing on the core semantics $S$ via the causal path ${Q}\rightarrow S\rightarrow A$, or on the irrelevant part $T$ via the causal path ${Q}\rightarrow T\rightarrow A$. Both pathways lead to language bias, e.g., the focus on the question in Figure 1 (a).
@@
The essence of correct reasoning lies in the model’s grasp of the underlying causal mechanisms necessary for solving a VQA problem. As illustrated in Figure 2), the green subgraph $G_{h}$ embodies this comprehension. It implies an understanding of how both the image and the question jointly influence the ground-truth result ${G}$ (via ${I}\rightarrow G$ and ${S}\rightarrow G$). Consequently, the model’s predictions should exhibit sensitivity and robustness to changes in the ground truth Stolfo et al. (2022), namely ${G}\rightarrow A$. No spurious correlations should directly affect $A$ unless it goes through the mediator $G$. Based on the above analyses, we articulate the concept of model sensitivity and robustness on VQA problems:
@@
  • Robustness, which assesses the undesired direct causal effect of unimodal biases, e.g., $I\rightarrow A$ and $T\rightarrow A$, where lower effects indicate greater robustness to input changes that do not alter the correct answer.
@@ 2.3 Causal Analysis of VQA Biases
    Direct intervention on the image $I$ by replacing an alternative image $I^{\prime}$.
@@
    Partially controllable interventions on $Q$. The question $Q$ can be modified in two ways: (i) altering both $S$ and $T$ , or (ii) altering $T$ but keeping $S$ unaffected.
@@
Next, we explain how we can obtain the causal effect from the interventions. Consider an intervention formulated as $\operatorname{do}\left(X:x\rightarrow x^{\prime}\right)$, where $X\in\{I,Q,T\}$ in the context of the VQA problem defined by $M=\{I,Q\}$. For simplicity, the probability distribution prior to the intervention, $\mathbb{P}(A\mid{I},{Q})$, is denoted by $P$, while the distribution following the intervention is represented as $P^{\prime}$. Following Pearl (1995), we measure the causal effect of each factor in our causal graph by employing the distance metric $\delta$ to gauge the divergence between distributions $P$ and $P^{\prime}$:
@@
It is important to note that $\operatorname{CE}$ denotes the causal effect, which can refer to either of the two components: the total causal effect ($\operatorname{TCE}$), signifying the joint effect across all causal paths from one variable to another, and the direct causal effect ($\operatorname{DCE}$), indicating the effect of the directed causal path devoid of intermediary variables Pearl (2022).
@@
Causal Effects of the Images
The causal effect size of the image on the answer can be obtained by conducting intervention on $I$, i.e.,
[…]
Note that, this $\operatorname{TCE}$ consist of two different paths of how $I$ affects $A$, as shown in Figure 2:
  • The desired decision-making route we aim for the model to adopt ( the ${I}\rightarrow G\rightarrow A$ path), where the model responds accordingly to alterations in the ground-truth answer.
  • The spurious correlation that the model might have learned (the ${I}\rightarrow A$ path), where the model depends on certain visual contexts potentially linked to their prevalence in the training corpus.
Maintaining $G$ constant during each intervention on $I$ allows us to quantify the $\operatorname{DCE}$ of $I$ on $A$, which refers to the strength of the path ${I}\rightarrow A$:
$\operatorname{DCE}({I}\rightarrow A):=\mathbb{E}_{{i}^{\prime}\sim\mathbb{P}({I}\mid G)}\left[\delta\left(P,P^{\prime}\right)\right],$ (6)
[…]
@@
As for the total causal effect of the question on the answer, we can quantify it by intervening on $Q$:
@@
It should be noted that the $\operatorname{DCE}({T}\rightarrow A$) equates to the $\operatorname{TCE}$ of $T$ on $A$, due to there is no mediator between $T$ and $A$. Given the impracticality of enumerating all possible perturbations of $T$, practical results within a specific subset of $T$ could be attained by intervening on $Q$ while leaving $S$ unaffected Stolfo et al. (2022). Besides, in the context of the VQA problem, solely intervening on $S$ without influencing the textual surface $T$ is unattainable. However, comprehending the causal impact of $S$ on $A$ becomes feasible through the comparison of two known quantities: $\operatorname{TCE}({Q}\text{ on }A)$ and $\operatorname{DCE}({T}\rightarrow A)$.
@@ 4.3 Causal Analysis of VQA Biases
Figrue 5 shows the distribution of options under the “Multi-choice” setting across various MLLMs. We observe that: 1) BLIP2 and GPT-4V often incorrectly choose options indicating language bias, aligning with our prior analysis on GPT-4V. 2) The proportion of either language or vision bias exceeded 40% across all the models, demonstrating the significant impact of unimodal biases on their predictions. 3) To some extent, models’ selection of semantically misleading options indicates some ability to combine visual and textual information, though not fully grasping the problem. This highlights the challenge our MORE dataset poses to current MLLMs. Please note that discrepancies may exist between the proportions of ground truth options presented here and the accuracy values reported in Table 2, as some models’ (e.g., mPLUG-Owl) outputs may not align with the provided options, affecting the count of valid answers.
Causal Effects of Images and Questions
@@
To further analyze the impact of vision bias and language bias on the model’s prediction, we assess the causal effects based on the definitions provided in Section 2.3. Specifically, we randomly select 100 samples for intervention and then measure the average of the effects over all instances to calculate $\operatorname{TCE}$ (corresponds to model’s sensitivity) and $\operatorname{DCE}$ (corresponds to models’ robustness). Overall, a higher $\operatorname{TCE}$ is desirable, indicating better sensitivity, while a lower $\operatorname{DCE}$ indicates better robustness. Implementation details of the intervention can be found in Appendix E.
@@ Appendix E Implementation Details of Intervention
In this section, we introduce the implementation details of interventions on images and questions.

### v3

@@ 2 A Causal Framework
Figure 2: Causal graph of MLLM’s Prediction on VQA problems. We use the green subgraph $G_{h}$ to represent the desired causal mechanisms and compare it with the undesired effects of unimodal biases. We quantify the causal effects of each factor by performing controlled interventions of the images ($I,E,C$) and of the questions ($Q,T$).
@@ 2.1 Problem Setup
We consider an entity-centric VQA problem, denoted by $M$, consisting of a question $Q$ and an image $I$. The image depicts a specific entity and the question is related to the entity. The question $Q:=(S,T)$ is composed of two distinct elements: the core semantic content ${S}$, which conveys the genuine intent of the question, and the textual surface form ${T}$, which is irrelevant to the question’s core meaning. The image $I:=(E,C)$, instead, contains the core entity of interest $E$ and the irrelevant visual context $C$. The model’s final answer/prediction is represented by $A$. In this paper, we use lowercase letters to represent an instance of its corresponding capital variable.
@@ 2.2 Causal Graph of MLLMs’ Prediction
${s}=f_{c_{1}}({q}),\ {e}=f_{c_{2}}({i}),\ g=f_{c_{3}}({s,e}),$ (1)
where the cognitive process $f_{c_{1}}$ is employed to decipher the core semantic meaning $s$ of the question $q$, and $f_{c_{2}}$ is employed to extract the core entity $e$ from the image $i$. Then, the function $f_{c_{3}}$ is utilized to combine $s$ and $e$, yielding the final result $g$. These mechanisms are depicted within the green subgraph $G_{h}$ of Figure 2.
In contrast, the process of a model to tackle the same VQA problem $m$ involves the following causal mechanisms:
$a=f_{b}({q},{i}),\vspace{-5pt}$ (2)
where $a$ represents the model’s prediction. The function $f_{\text{b}}$ operates as a black box, leaving uncertainty regarding which aspects of $q$ and $i$ are considered by the model and how they interact with each other.
To further analyze how a model tackles the VQA problem, we draw all possible causal mechanisms that might take place in the complete causal graph in Figure 2. Notable causal mechanisms include:
Language Bias
The model may directly process the question $Q$ in two ways: by focusing on the core semantics $S$ via the causal path ${Q}\rightarrow S\rightarrow A$, or on the irrelevant part $T$ via the causal path ${Q}\rightarrow T\rightarrow A$. Both pathways lead to language bias, e.g., the focus on “the next World Cup ” in Figure 1 (a).
@@
The model may attend directly to the entity $E$ of the image $I$ via $I\rightarrow{E}\rightarrow A$, or to the irrelevant part $C$ via $I\rightarrow{C}\rightarrow A$. Both pathways lead to the emergence of vision bias, e.g., the focus on the “The Shard” entity in Figure 1 (b).
@@
The essence of correct reasoning lies in the model’s grasp of the underlying causal mechanisms necessary for solving a VQA problem. As illustrated in Figure 2), the green subgraph $G_{h}$ embodies this comprehension. It implies an understanding of how both the image and the question jointly influence the ground-truth result ${G}$ (via ${E}\rightarrow G$ and ${S}\rightarrow G$). Consequently, the model’s predictions should exhibit sensitivity and robustness to changes in the ground truth Stolfo et al. (2022), namely ${G}\rightarrow A$. No spurious correlations should directly affect $A$ unless it goes through the mediator $G$. Based on the above analyses, we articulate the concept of model sensitivity and robustness on VQA problems:
@@
  • Robustness, which assesses the undesired direct causal effect of unimodal biases, e.g., $C\rightarrow A$ and $T\rightarrow A$, where lower effects indicate greater robustness to input changes that do not alter the correct answer.
@@ 2.3 Causal Analysis of VQA Biases
    Interventions on $Q$. The question $Q$ can be modified in two ways: (i) altering both $S$ and $T$ , or (ii) altering $T$ but keeping $S$ unaffected.
@@
    Interventions on $I$. The image $I$ can be replaced with an alternative image $I^{\prime}$ in three ways: (i) altering both $E$ and $C$ , or (ii) altering $C$ but keeping $E$ unaffected, or (iii) altering $E$ but keeping $C$ unaffected.
Note that we do not solely alter $S$ within $Q$, because it is not feasible to intervene on the core semantics $S$ of a question without affecting the surface text $T$ of it.
@@
Next, we explain how we can obtain the causal effect from the interventions. Consider an intervention formulated as $\operatorname{do}\left(X:x\rightarrow x^{\prime}\right)$, where $X\in\{I,E,C,Q,T\}$ is a factor in the context of the VQA problem defined by $M=\{I,Q\}$. For simplicity, the probability distribution prior to the intervention, $\mathbb{P}(A\mid{I},{Q})$, is denoted by $P$, while the distribution following the intervention is represented as $P^{\prime}$. Following Pearl (1995), we measure the causal effect of each factor in our causal graph by employing the distance metric $\delta$ to gauge the divergence between distributions $P$ and $P^{\prime}$:
@@
It is important to note that $\operatorname{CE}$ denotes the causal effect, which can refer to either of the two forms: the total causal effect ($\operatorname{TCE}$), signifying the joint effect across all causal paths from one variable to another; and the direct causal effect ($\operatorname{DCE}$), indicating the effect of the directed causal path devoid of intermediary variables Pearl (2022).
@@
For the total causal effect of the question on the answer, we can quantify it by intervening on $Q$:
@@
Note that, this $\operatorname{TCE}$ contains two different types of paths that show how $Q$ affects $A$, as illustrated in Figure 2:
  • The desired decision-making route we aim for the model to adopt ( the ${Q}\rightarrow S\rightarrow G\rightarrow A$ path), where the model responds accordingly to alterations in the ground truth answer.
  • The spurious correlation that the model might have learned (e.g., the ${Q}\rightarrow T\rightarrow A$ path, where the model depends on certain linguistic contexts potentially linked to their prevalence in the training corpus).
@@
It should be noted that the $\operatorname{DCE}({T}\rightarrow A$) equates to the $\operatorname{TCE}$ of $T$ on $A$, due to there is no mediator between $T$ and $A$. Besides, as discussed in “Causal Interventions”, solely intervening on $S$ without influencing the textual surface $T$ is unattainable. However, comprehending the causal impact of $S$ on $A$ becomes feasible through the comparison of two known quantities: $\operatorname{TCE}({Q}\text{ on }A)$ and $\operatorname{DCE}({T}\rightarrow A)$.
Causal Effects of the Images
It is evident from Figure 2 that the causal structure of images is exactly the same as that of questions. Therefore, the total causal effect of the image $I$ on the answer can be directly obtained in a similar manner, i.e.,
[…]
Likewise, maintaining $E$ constant during each intervention on $I$ allows us to quantify the $\operatorname{DCE}$ of the irrelevant visual context $C$ on $A$, which refers to the strength of the path ${C}\rightarrow A$:
$\operatorname{DCE}({C}\rightarrow A):=\mathbb{E}_{{i}^{\prime}\sim\mathbb{P}({I}\mid E)}\left[\delta\left(P,P^{\prime}\right)\right],$ (8)
[…]
Maintaining $C$ and $G$ constant during each intervention on $I$ instead allows us to quantify $\operatorname{DCE}({E}\rightarrow A)$:
$\operatorname{DCE}({E}\rightarrow A):=\mathbb{E}_{{i}^{\prime}\sim\mathbb{P}({I}\mid C,G)}\left[\delta\left(P,P^{\prime}\right)\right],$ (9)
[…]
Note that in practical situations, it is significantly challenging to intervene on the core entity $E$ while simutaniously controlling for the irrelevant visual context $C$, the ground truth answer $G$, and the question $Q$. Therefore, this type of intervention will not be considered in subsequent experiments.
@@ 4.3 Causal Analysis of VQA Biases
In Figrue 5, we show the distribution of options under the “Multi-choice” setting across various MLLMs. We observe that: 1) BLIP2 and GPT-4V often incorrectly choose options indicating language bias, aligning with our prior analysis on GPT-4V. 2) The proportion of either language or vision bias exceeded 40% across all the models, demonstrating the significant impact of unimodal biases on their predictions. 3) To some extent, models’ selection of semantically misleading options indicates some ability to combine visual and textual information, though not fully grasping the problem. This highlights the challenge our MORE dataset poses to current MLLMs. Please note that discrepancies may exist between the proportions of ground truth options presented here and the accuracy values reported in Table 2, as some models’ outputs (e.g., mPLUG-Owl) may not align with the provided options, thus affecting the count of valid answers.
@@
Causal Effects of Images and Questions
To further analyze the impact of vision bias and language bias on the model’s prediction, we assess the causal effects based on the definitions provided in Section 2.3. Specifically, we randomly select 100 samples for intervention and then measure the average of the effects over all instances to calculate $\operatorname{TCE}$ (corresponds to model’s sensitivity) and $\operatorname{DCE}$ (corresponds to models’ robustness). We implement two straightforward and natural intervention ways separately for questions and images. For questions: (i) we change both the core semantics and the textual form, or (ii) we only change the textual form while controlling to keep the core semantics unchanged. For images: (i) we change both the entities of interest and the visual context, or (ii) we only change the visual context while keeping the core entities unchanged. As discussed below Equation (9), we do not conduct sole interventions on the core entity while concurrently controlling for the irrelevant visual context, the ground truth answer, and the question. This is because it is extremely difficult to find such qualified samples to calculate $\operatorname{DCE}({E}\rightarrow A)$ in practical scenarios. Overall, a higher $\operatorname{TCE}$ is desirable, indicating better sensitivity, while a lower $\operatorname{DCE}$ indicates better robustness. Implementation details of the intervention can be found in Appendix E.
@@ Appendix E Implementation Details of Intervention
In this section, we introduce the implementation details of interventions on questions and images. Given the impracticality of enumerating all possible perturbations of $T$ and $R$, practical results within a specific subset of $T$ and $R$ could be attained by intervening on $Q$ while leaving $S$ unaffected, and intervening on $I$ while leaving $E$ unaffected, respectively Stolfo et al. (2022).